OpenStudy (anonymous):

7/8 plus 1/2? I don't get this at ALL. I think I must have been sleeping when my teacher talked about it.

OpenStudy (anonymous):

1/2 = 4/8. 7/8 + 4/8 = 11/8 = 1 3/8 When the denominators are the same, you can add the nominators.

OpenStudy (anonymous):

7/8 + 1/2 = 7/8 + 4/8 " cuz we multiply by 4 " 7/8 + 4/8 = 11/ 8

OpenStudy (anonymous):

Oh, Thank you! This site is SO cool!